Income Statement
Revenue
Total revenue $470,100
Cost of Revenue $149,700
Gross Profit $320,100
Operating Exp
Selling General & Admin $57,800
other $17,400
Total Operating Exp $269,900
Operating Income or Loss $200,200
Income from Continuing Operations
Expenses Net/Total other Income $114,400
Earnings before Int & Tax $200,200
Interest Exp -$32,200
Income before Tax $314,600
Income Tax Exp $100
Minority Interest $23,300
Net Income from Continuing Ops $314,500
Net Income   $295,300
Net Income applicable to common shares $230,700
Balance Sheet
Current Assets
Cash and Cash Equ $137,700
Net Receivables $41,000
Inventory $12,400
Total current assets $203,600
Long term investments $424,700
Property Plant & Equipment $736,500
Other Assets $1,700
Total Assets $1,366,500
Current Liabilities
Accounts Payable $15,600
Other Current Liabilities $19,400
Total Current Liabilities $54,900
Long term debt $1,844,000
Other Liabilities $9,200
Minority Interest $23,300
Total Liabilities $1,932,400
Stockholders' Equity
Common Stock -$589,200
Total Stockholder Equity -$589,200
Net Tangible assets -$589,200

Based on the income statement and balance sheet above, answer the following (show work and formulas please)

1. Current Ratio = CA/CL =

2. Inventory Turnover

3. Total asset utilization (turnover)

4. Debt Ratio = Total debt (Liabilities)/ Total Assets

5. Debt to Equity ratio

6.  Basic earning power

7. ROE
Answer #1
Current ratio      3.71 times Current assets/Current liabilities =203600/54900
Inventory turnover    37.91 times Sales/Inventory =470100/12400
Total assets turnover      0.34 times Sales/Total assets =470100/1366500
Debt ratio      1.41 times Total debt/Total assets =1932400/1366500